# Hyper-STTN: Social Group-aware Spatial-Temporal Transformer Network for Human Trajectory Prediction with Hypergraph Reasoning

### v1

@@ IV-A1 Dataset
We have conducted experiments on 5 commonly used public pedestrian trajectory datasets collected from real-world scenarios. As shown in Table I, the training and validation datasets consist of the ETH, HOTEL, ZARA1, ZARA2, and UNIV subsections. These datasets record the temporal location information of pedestrians in world coordinates, each identified by an individual index. Our experiments used a leave-one-out cross-validation evaluation strategy, where the last one scenario is reserved for testing in each experiment. The datasets are also augmented through various operations such as shifting and rotating to enhance the training process.
@@ IV-A2 Evaluation Metrics
[…]
@@ IV-B1 Baselines
We compared our model with several existing SOTA algorithms. The algorithms included in the comparison are Social-LSTM [5], Social-Attention [14], Social-GAN [7], Trajectron++ [10], STAR [8], AgentFormer [9], Social-Implicit [28], GroupNet [6], and EqMotion [29]. We trained and tested all baseline networks using the same training datasets and validation datasets obtained from public sources, and we followed the parameters and configurations outlined in their respective original papers.
@@ IV-B2 Ablation Models
We designed three ablation models to investigate the influence of different subsection blocks on performance. First, STTN is designed to infer underlying pedestrian interactions based soley on the spatial-temporal transformer, without the hypergraph network block from Hyper-STTN. Meanwhile, HGNN is configured to model crowded motion dependencies using a set of multiscale hypergraph constructions and hypergraph convolution operations. The final ablation model (Hyper-STTN+MLP) removes the multimodal transformer block and employs a MLP (multilayer perceptron) for group-wise and pair-wise feature fusion. Additionally, the CVAE structure serves as a decoder for each ablation model. And training configurations, such as the training dataset and initialized parameters, are shared among the ablations models and Hyper-STTN.
[…]
@@ IV-B3 Training Details
To train the model, we use the Adam optimizer [30] with a learning rate $1\times 10^{-4}$ and a decay rate (50% per 100 epochs). Additionally, Gaussian noise is added to enforce the robustness of the network. For more parameters and details, please refer to our project website.
@@
TABLE II: The results of ADE/FDE on ETH-UCY in deterministic tests.
Deterministic ADE / FDE S-LSTM[5] CVPR16 Trajectron++[10] ECCV20 STAR[8] ECCV20 EqMotion[29] CVPR23 Hyper-STTN (ours)
ETH 1.09/2.35 1.02/2.00 0.56/1.11 0.96/1.92 0.52/1.02
HOTEL 0.79/1.76 0.33/0.62 0.26/0.50 0.35/0.67 0.22/0.41
UNIV 0.67/1.40 0.53/1.19 0.52/1.15 0.50/1.10 0.50/1.16
ZARA1 0.47/1.00 0.44/0.99 0.41/0.90 0.39/0.86 0.41/0.88
ZARA2 0.56/1.17 0.32/0.73 0.31/0.71 0.30/0.68 0.32/0.72
AVG 0.72/1.54 0.53/1.11 0.41/0.87 0.49/1.03 0.39/0.84
@@ IV-C Qualitative Results
We conducted two trajectory prediction tasks respectively, wherein predictors only generate single path in deterministic conditions, and output 20 potential paths to evaluate with the best one of such 20 samples in stochastic conditions. To adapt to deterministic situations, we modified the CVAE decoder of our model to use a MLP, allowing us to omit the second term $f_{\rm KL}()$ in $\mathcal{L}_{dis}$ loss and the entire $\mathcal{L}_{enc}$ loss.
[…]
However, despite transformer demonstrates the amazing effectiveness on pair-wise interaction abstraction, the overlook of group-wise interaction limits the further performance improvement, particularly in dense scenarios individual motion is more affected by crowded groups. Such as GroupNet [6] ($0.26/0.49$) is better than STAR [8] ($0.31/0.62$) on UNIV dataset that owns more dense crowded scenarios. Therefore, herein we leverage the ability of hypergraph construction and convolution on group-wise interaction inference and transformer on pair-wise feature understanding to address crowded movement strategies. Apart from that, the ablation model HGNN exhibits a similar performance with GroupNet on several datasets.
(a) Sample-1
(b) Sample-1
(c) Sample-2
(d) Sample-2
Figure 5: Comparison of trajectories visualizations: The trajectories visualized for HyperSTTN and other algorithms tested on the same test case.
@@ IV-D1 Effect of Hypergraph Neural Network
We address three components of Hyper-STTN, which include the hypergraph neural network, spatial-temporal transformer, and multi-modal transformer. As shown in Table I, HGNN performs approximately as well as GroupNet on many datasets. In general, hypergraph construction can effectively enhance network performance as the number of pedestrians increases. For example, in Fig. 5d, the ablation model STTN exhibits serious forecasting errors of many agents, both in terms of location and orientation, when compared to HGNN. This is attributed to STTN’s lack of consideration for group-wise features.
@@ IV-D3 Effect of Multi-Modal Transformer
Due to the heterogeneity of group-wise and pair-wise embeddings, Hyper-STTN leverages a multi-modal transformer to merge them, capturing cross-modal attention to highlight meaningful correlations among modalities. In contrast, the limited fusion mechanism in the ablation model Hyper-STTN+MLP underperforms Hyper-STTN in every dataset, as shown in Table I. This is because the direct concatenation operation in the MLP overlooks potential relationships among cross-modal features.
@@
(c) Crowd Group (scale=2)
(d) Crowd Group (Scale=3)
Figure 6: The illustration of group-wise and pair-wise interactions in Hyper-STTN: In attention maps from the spatial temporal transformer, the red circle represents the attention score of current agent, and the relative importance of neighboring agents is captured by the radius of black circles. Additionally, crowd group-wise interactions are described by hypergraph constructions with different scales in subfigures (c) and (d).
@@ V Conclusion
In this work, we introduced Hyper-STTN, a hypergraph-based hybrid spatial-temporal transformer for improved pedestrian trajectory prediction. Hyper-STTN captures both group-wise and pair-wise interactions in crowd dynamics using hypergraphs and transformer blocks. It has outperformed the state-of-the-art prediction algorithm on the ETH-UCY dataset. Future work could explore further enhancements to Hyper-STTN, including refining its ability to handle complex scenarios and scaling it for real-time applications.

### v2

@@ IV-A1 Dataset
We have conducted experiments on several commonly used public pedestrian trajectory datasets collected from real-world scenarios. The training and validation datasets are consisted of the ETH-UCY [28], NBA, and SDD [29] datasets. These datasets record the temporal location information of pedestrians in world coordinates, each identified by an individual index. Additionally, the experiments of ETH-UCY are conducted on the leave-one-out cross-validation evaluation strategy, where the last one subsection is reserved for testing in each experiment.
@@ IV-A2 Evaluation Metrics
In our experiments, we employed the evaluation metrics Average Displacement Error (ADE) and Final Displacement Error (FDE) [4] to assess prediction accuracy. The ADE/FDE and ADE20/FDE20 metrics measure prediction performance by calculating the Euclidean distance between the prediction and ground truth, either on average or at a single endpoint.
@@ IV-B1 Baselines
We compared our model with several existing SOTA algorithms. The algorithms included in the comparison are Social-Attention [13], Social-GAN [6], Social-STGCNN [15], Trajectron++ [9], STAR [7], PECNet [30], AgentFormer [8], Social-Implicit [31], GroupNet [5], and EqMotion [32]. We trained and tested all baseline networks using the same training datasets and validation datasets obtained from public sources, and we followed the parameters and configurations outlined in their original papers.
@@ IV-B2 Ablation Models
We designed three ablation models to investigate the influence of different subsection blocks on performance. First, STTN is designed to infer underlying pedestrian interactions based solely on the spatial-temporal transformer, without the hypergraph network block from Hyper-STTN. Meanwhile, HGNN is configured to model crowded motion dependencies using a set of multiscale hypergraph constructions and hypergraph convolution operations. The final ablation model (Hyper-STTN+MLP) removes the multimodal transformer block and employs a MLP (multilayer perceptron) for group-wise and pair-wise feature fusion. Additionally, the CVAE structure serves as a decoder for each ablation model. And training configurations, such as the training dataset and initialized parameters, are shared among the ablations models and Hyper-STTN.
TABLE I: The crowded trajectory forecasting best-of-20 stochastic sampled results of minADE20 / minFDE20 on ETH-UCY dataset.
Stochastic ADE20 / FDE20 Social- Attention [13] ICRA18 Social- GAN [6] CVPR18 Trajectron ++ [9] ECCV20 STAR [7] ECCV20 Agent- Former [8] ICCV21 Social- Implicit [31] ECCV22 GroupNet [5] CVPR22 EqMotion [32] CVPR23 STTN (ablation) HGNN (ablation) Hyper-STTN +MLP (ablation) Hyper-STTN (ours)
@@
TABLE II: The crowded trajectory forecasting best-of-20 stochastic sampled results of minADE20 / minFDE20 on NBA &\&& SDD dataset.
Stochastic ADE20 / FDE20 Social-GAN [6] CVPR18 Social-STGCNN [15] CVPR20 Trajectron++ [9] ECCV20 PECNet [30] ECCV20 GroupNet [5] CVPR22 STTN (ablation) HGNN (ablation) Hyper-STTN+MLP (ablation) Hyper-STTN (ours)
NBA-1s 0.41/0.62 0.34/0.48 0.30/0.38 0.35/0.58 0.34/0.48 0.39/0.58 0.37/0.59 0.35/0.48 0.30/0.37
NBA-2s 0.81/1.32 0.71/0.94 0.59/0.82 0.68/1.23 0.62/0.95 0.72/1.05 0.64/0.95 0.65/0.89 0.58/0.81
NBA-3s 1.19/1.94 1.09/1.77 0.85/1.24 1.01/1.76 0.87/1.31 1.07/1.51 0.98/1.44 0.95/1.51 0.84/1.21
NBA-4s 1.59/2.41 1.53/2.26 1.15/1.57 1.31/1.79 1.13/1.69 1.24/1.99 1.18/1.89 1.21/1.75 1.01/1.52
SDD 27.23/41.44 20.8/33.2 19.30/32.70 9.96/15.88 9.31/16.11 11.17/19.28 11.13/18.43 10.11/18.41 9.05/15.14
@@ IV-B3 Training Details
To train the model, we use the Adam optimizer with a learning rate $1\times 10^{-4}$ and a decay rate (50% per 100 epochs). Additionally, Gaussian noise is added to enforce the robustness of the network. For more parameters and details, please refer to our project website.
@@ IV-C Qualitative Results
We conducted the trajectory prediction task with respect to stochastic conditions, which generate 20 potential paths to evaluate with the best one of total samples. As shown in Tables I and II, Hyper-STTN is compared with several baselines and three ablation models on the ETH-UCY dataset, NBA dataset and SDD dataset. Hyper-STTN demonstrates the best performance among current SOTA frameworks, both in ETH-UCY, NBA, and SDD configurations. For instance, in the ETH-UCY dataset, Hyper-STTN not only enhances the $12.5\%$ ADE20 and $3.2\%$ FDE20 from EqMotion’s [32] 0.40/0.61 to 0.35/0.59 on ETH dataset, but also improves the average $14.7\%/6.5\%/3.4\%/10.6\%$ ADE20 and $22.9\%/14.7\%/7.6\%/10.1\%$ FDE20 performance from Groupnet’s [5] on NBA dataset. Highlighting the effectiveness of Hyper-STTN in crowd latent interaction inference and trajectory forecasting.
The insights of deep learning-inspired trajectory prediction framework have motivated the research paradigm that implicitly capture latent social interaction among spatial and temporal dimensions to estimate pedestrian potential movements. Whereas, the deployments of LSTM [4], Attention [13], to transformer [7] have increased the average performance of pedestrians’ trajectory predictors with respect to the more accurate high-level feature inference on advanced framework. Thus, the transformer-based backbones are leveraged by Hyper-STTN to address fundamental social interactions and temporal dependencies among crowded movements. More specifically, the ablation model STTN, which is solely driven by a transformer like STAR. From our experiments, we observed that the aforementioned enrichment is a result of the introduction of masked attention and cross-modal attention mechanisms. These mechanisms not only mitigate the influence of agents flicker when compared to vanilla self-attention in the fixed pre-defined transformer data scale formulation but also enhance the adaptability for spatial-temporal fusion.
However, despite transformer demonstrates the amazing effectiveness on pair-wise interaction abstraction, the overlook of group-wise interaction limits the further performance enhancements, particularly in dense scenarios individual motion is more sensitive for the movements of crowded groups. Such as GroupNet [5] ($0.26/0.49$) is better than STAR [7] ($0.31/0.62$) on UNIV dataset that owns more dense crowded scenarios. Therefore, herein we leverage hypergraph construction on group-wise interaction inference and transformer on pair-wise feature understanding to address crowded movement strategies. Apart from that, the ablation model HGNN exhibits a similar performance with GroupNet on several datasets.
(a) Instance-1
(b) Instance-1
(c) Instance-2
(d) Instance-2
Figure 5: Comparison of Trajectories Visualizations: The trajectories visualized for Hyper-STTN and other algorithms tested on the same test case.
@@ IV-D1 Effect of Hypergraph Neural Network
We address three components of Hyper-STTN, which include the hypergraph neural network, spatial-temporal transformer, and multi-modal transformer. As shown in Table I, HGNN performs approximately as well as GroupNet on many datasets. In general, hypergraph construction can effectively enhance network performance as the number of pedestrians increases. For example, in Fig. 5d, the ablation model STTN exhibits serious forecasting errors of many agents, both in terms of location and orientation, when compared to HGNN. And the crowd hypergraph group constructions are shown in in Fig. 6 which is attributed to STTN’s lack of consideration for group-wise features.
@@ IV-D3 Effect of Multi-Modal Transformer
Due to the heterogeneity of group-wise and pair-wise embeddings, Hyper-STTN leverages a multi-modal transformer to merge them, capturing cross-modal attention to highlight meaningful correlations among modalities. In contrast, the limited fusion mechanism in the ablation model Hyper-STTN+MLP underperforms Hyper-STTN in every dataset, as shown in Table I and Table II. This is because the direct concatenation operation in the MLP overlooks potential relationships among cross-modal features.
@@
(c) Crowd Group (scale=3)
(d) Crowd Group (Scale=4)
Figure 6: The illustration of group-wise and pair-wise interactions in Hyper-STTN: The pair-wise attention maps are illustrated on subfigures (a) and (b), the red and black circles represent the attention score of current agent and neighboring agents respectively. Additionally, crowd group-wise interactions are described by hypergraphs with different scales in subfigures (c) and (d).
@@ V Conclusion
In this work, we introduced Hyper-STTN, a hypergraph-based hybrid spatial-temporal transformer for trajectory prediction tasks. Hyper-STTN captures both group-wise and pair-wise interactions in crowd dynamics using hypergraphs and transformer blocks. It has outperformed the state-of-the-art prediction algorithm on the ETH-UCY, NBA, and SDD dataset. Future work could explore further enhancements to Hyper-STTN, including refining its ability to handle complex scenarios and scaling it for real-time applications.